OpenStudy (anonymous):

1. Solve the equation 2x^2+10x=0 . Check your solution(s) and state the final solution set. Show all your work.

OpenStudy (anonymous):

\[2x^2+10x=0\] put x in evidence \[x*(2x+10)=0\] so \[x=0~~and~~2x+10=0\] \[2x+10=0\] \[2x=-10\] \[x=-5\] \[\therefore~x=0~~and~x=-5\]